An ice cream container is cylindrical with diameter 30cm and height 50cm and is filled with ice
cream. The ice cream scoop makes ice cream scoops of diameter 6cm. How many one scoop
cones can be made from one cylinder of ice cream?

Answer by FrankM(1040)

Your cylinder has a base radius of 15cm, πr^2 is the base area, or 225π cm^2.
Multiply by the height to get volume 225π * 50 = 11,250π cm^3.
A sphere has v=4/3πr^3 or 4/3∏(3)^3 = 36π cm^3
The π cancels, and we are left with 11,250/36 = 312.5 scoops.